What is the ratio of acetate ion to acetic acid (Ka = 1.76×10–5) in a solution containing these compounds at pH 3.17? Write your answer as a decimal and not as a fraction.
Answer #1

pKa = - log (Ka)

= - log(1.76*10^-5)

= 4.7545

pH = pKa + log (CH3COO- / CH3COOH)

CH3COO- / CH3COOH = 10^(pH - pKa)

= 10^(3.17 - 4.7545)

= 0.026

any query please comment

if satisfied please rate it thanks